### v1

# Underwater Image Enhancement by Diffusion Model with Customized CLIP-Classifier

Abstract
In this paper, we propose a novel underwater image enhancement method, by utilizing the multi-guided diffusion model for iterative enhancement. Unlike other image enhancement tasks, underwater images suffer from the unavailability of real reference images. Although existing works exploit synthetic images, manually selected well-enhanced images as reference images, to train enhancement networks, their enhancement performance always comes with subjective preferences that are inherited from the manual selection. To address this issue, we also use the image synthesis strategy, but the synthetic images derive from in-air natural images degraded into corresponding underwater images, guided by the underwater domain. Based on this strategy, the diffusion model can learn the prior knowledge of image enhancement from the underwater degradation domain to the real in-air natural domain. However, it is inevitable to fine-tune the model to suit downstream tasks, and this may erase the prior knowledge. To mitigate this, we combine the prior knowledge from the in-air natural domain with Contrastive Language-Image Pretraining (CLIP) to train a classifier for controlling the diffusion model generation process. Moreover, for image enhancement tasks, we find that the image-to-image diffusion model and the CLIP-Classifier mainly act in the high-frequency region during the fine-tuning process. Therefore, we propose a fast fine-tuning strategy focusing on the high-frequency region, which can be up to 10 times faster than the traditional strategy. Extensive experiments demonstrate that our method, abbreviated as CLIP-UIE, exhibit a more natural appearance. The source code and pre-trained models are available on the project homepage: https://github.com/OUCVisionGroup/CLIP-UIE.
Index Terms:
Underwater image enhancement, Diffusion model, Customize prompt, CLIP, Fine-tuning strategy
I Introduction
Underwater image enhancement (UIE) is the process of generating a visually clear image that is consistent with an input-degraded underwater image, which often exhibits color casts, low contrast, and poor visibility [1, 2, 3]. It plays a significant role in vision-based research for marine and underwater tasks, e.g. marine animal detection [4], ecosystem protection [5], and underwater robot application [6], etc.
By reviewing typical UIE strategies, early approaches, based on the physical model, focus on the accurate estimation of medium transmission that represents the percentage of the scene radiance reaching the camera [7, 8, 9]. Though physical model-based methods can achieve promising performance in some cases, they tend to produce unnatural and poor results in challenging underwater scenarios [10]. Recent approaches based on deep learning technology have gradually outperformed traditional physical model-based approaches. [11, 12, 13]. While simple regression-based methods with feedback convolutional networks may work for image enhancement in simple underwater scenarios, they often lacks generalisation and robustness for complex scenes as they cannot fully model the real-underwater degradation. The image-to-image diffusion models have been successful in learning complex empirical distributions of images, enabling the transformation of images between different domains through iterative refinement steps.
However, image-to-image diffusion models heavily rely on paired datasets, and the inability to capture clear real-world reference images poses a challenge for UIE [14, 15, 16]. To address this issue, researchers employed various approaches to synthesize multiple enhanced images of a single image, and then manually select the visually optimal one as reference [17, 18]. Nevertheless, the synthesized paired datasets contain subjective preferences, reflected in the enhanced results, resulting in natural domain shift and weakened generalization. As for the unsupervised or semi-surprised strategies, they can also generate pleasing underwater images [19, 18], without paired datasets or relying on limited paired samples, but the performance is not superior to the fully supervised method. The other alternative approach is to combine a pre-trained diffusion model, which captures the prior knowledge of the mapping transitions between different image domains, with custom datasets, thus reducing dependency on paired datasets [20]. However, the fine-tuning, to directly continue training with the additional paired data, may destroy the valuable prior knowledge [21]. Moreover, another limitation, the quality of the enhancement result is limited by the reference domain due to the lack of real image guidance, resulting in unnatural and unrealistic enhanced images.
To address the above limitations, we propose CLIP-UIE, a new approach to treat UIE as a conditional underwater image generation, inspired by recent works on Denoising Diffusion Probabilistic Models [22, 23, 24, 25] and CLIP guidance for image manipulation [26, 27, 28]. The proposed method is trained with a denoising objective to iterative refinement enhancement images by giving a degraded image. Initially, the key is to obtain a large number of underwater paired datasets. We utilize the image synthesis strategy based on color transfer [29, 30] to degrade in-air natural images into the real underwater domain, synthesising the corresponding degraded underwater images, which is free of any subjective preferences. Subsequently, we train the basic image-to-image diffusion model on this synthetic paired dataset to learn the prior knowledge of the mapping transitions from the underwater degradation domain to real in-air natural domain. This prior knowledge can adverse impact of unrealistic and unnatural synthetics effectively. However, it is inevitable to fine-tune the model to suit downstream tasks, and this may erase the prior knowledge, leading to overfitting and mode collapse [21]. Song et al. [31] and Jonathan et al. [32] find that a pre-trained diffusion model can be conditioned using the gradients of a classifier. Thus, we can exploit a trained classifier with prior knowledge, combined with custom datasets to jointly fine-tune the diffusion model to avoid overfitting and mode collapse, bridging the domain gap between synthetic and real images. It has been shown that Vision-Language models encapsulate generic information, that can be paired with generative model to provide a simple and intuitive text-driven interface for image generation and manipulation [33], particularly CLIP. Although CLIP can serve as a classifier to distinguish in-air natural and underwater degraded images to a certain extent, using its gradients directly to control the reverse diffusion process is still non-trivial. Moreover, it is barely possible to find accurate prompts to describe the types of underwater degradation, and rewriting similar prompts easily leads to huge differences in CLIP score, as shown in Fig. 1. We train a CLIP-Classifier that contains prior knowledge of the in-air natural domain by constraining text-image similarity in the CLIP embedding space. Moreover, for image enhancement tasks, we find that the image-to-image diffusion model and the CLIP-Classifier mainly act in the high-frequency region during the fine-tuning process. Therefore, we propose a fast fine-tuning strategy that focuses on the high-frequency range and can be up to 10 times faster. The contributions of this paper are summarized as follows:
  • We propose to use color transfer to degrade 500k in-air natural images from INaturalist [34] into underwater images guided by the real underwater domain, overcoming subjective preferences introduced by manual selection of reference images. The pre-trained image-to-image diffusion model is trained from scratch on this synthetic dataset.
  • We propose a CLIP-Classifier that inherits prior knowledge of the in-air natural domain, and then combine it with custom datasets to jointly fine-tune the diffusion model to mitigate catastrophic forgetting and mode collapse. Experiments and ablation studies verify the good performance of the proposed CLIP-UIE and it breaks the limitations of the reference domain to a certain extent.
  • We find that for image enhancement tasks, which require consistent content, the image-to-image diffusion model and the CLIP-Classifier mainly act in high-frequency regions. Therefore, we propose a new fine-tuning strategy that acts only on high-frequency regions, which significantly improves the fine-tuning speed, even up to 10 times.
Figure 1: CLIP with proper prompts can serve as a classifier to evaluate the image quality and distinguish between in-air natural and underwater images. However, our simple approximate rewriting of the prompt results in a dramatic change in CLIP score. On the contrary, our CLIP-Classifier with learned prompt shows more robust results for different types of in-air natural/underwater scenes.
II Related Works
II-A Underwater Image Enhancement
Several approaches have been proposed in the underwater enhancement image, categorizing into two groups: traditional methods and deep learning-based methods. For the traditional methods, researchers focus on directly adjusting image pixel values to produce a subjectively and visually appealing image, using techniques such as image fusion [35], histogram equalization [36], and pixel distribution [37], etc. Ancuti et al. [38] propose the fusion-based underwater image enhancement method, exploiting two images derived from a color-compensated and white-balanced version to prompt color contrast. However, these traditional methods are not sensitive to the types of underwater degradation, due to the omission of the underwater imaging mechanism. To improve the visual quality, physical model-based methods directly estimate the parameters of underwater imaging models based on prior information. For example, Fu et al. [7] utilize the retinex-based approach to capture a clean underwater image by combining enhanced reflectance and illumination. To further promote the performance, Zhuang et al. [9] improve the retinex-based algorithm by introducing Bayesian to estimate underwater medium transmissions. Nevertheless, owing to the scarcity of prior information, the enhanced image easily becomes unnatural artifacts and over/under-enhanced.
Recently, deep learning has made significant advances in computer vision and image processing, there are serial attempts made to improve the performance of underwater image enhancement by data-driven strategy. However, underwater images suffer from the unavailability of real reference images, leading researchers to synthesize paired datasets. Concretely, Li et al. [17] construct an underwater image benchmark (UIEB) by selecting the best result of various approaches as the reference. Nevertheless, the synthetic paired datasets cannot avoid subjective preferences from the manual selection. To relax the requirement of paired underwater images, several GAN-based weakly supervised methods can generate pleasing underwater images [19, 18], but the performance is not superior to the fully supervised method.
The strategy of using in-air natural images to synthesise paired underwater degradation datasets provides a good solution. Du et al. [20] transform in-air natural images into the pseudo underwater image domain by Koschmieder’s light scattering model, utilizing the prior knowledge of in-air natural images to counteract the adverse impact of unreal synthetic. However, the pseudo-underwater domain based on the light scattering model cannot fully model the complexity of real-world underwater degradation. To solve this issue, we propose a new synthesis strategy based on color transfer [29] to yield paired datasets, degrading in-air natural images to real underwater domain. By this means, we introduce more prior knowledge: the mapping between the real underwater degradation domain and the real in-air natural domain.
II-B Diffusion Models
Diffusion models recently have impressive performance on image generation by inverting image noising process with reverse-time stochastic differential equations (SDEs) [39]. However, this inversion process tends to change the image context. To condition the denoising process, Classifier Guidance [32, 31] is an effective method to improve the sample quality of conditional diffusion models with the gradients of the classifier. An alternative strategy is to establish the relationship between input and conditional information in the model’s latent space [40, 25]. Controlling diffusion models facilitates personalisation, customization, or task-specific image generation. Hence, the diffusion models have been successfully applied to various image enhancement tasks. The image-to-image diffusion model was first introduced by Saharia et al. [25] and has been exploited for image enhancement [41, 20], inpainting [42], and super-resolution [25], etc.
Conditional diffusion models can learn the mapping between different image domains [21]. Based on this observation, our proposal uses the aforementioned color transfer to synthetic images, and trains the conditional diffusion model to learn the mapping between the real underwater degradation domain and the real in-air natural domain. However, this model will be guided by the latter condition without real-underwater reference, and this may destroy prior knowledge. To mitigate this issue, we train a classifier to discriminate between in-air natural and underwater degraded images, integrating the prior knowledge of the in-air natural domain into this classifier, and then combine it with custom datasets to jointly fine-tune the diffusion model, enabling the transfer of enhancement results to both the natural and reference domains.
II-C CLIP-based Image Manipulation
The CLIP model benefits from 400 million curated image-text pairs can predict the correct pairings of a batch of (image, text) unseen examples, and consists of a text encoder and an image encoder by jointly training [43]. It has shown remarkable performance in text-to-image generation. Gal et al. [33] and Patashnik et al. [44] use CLIP models to modify StyleGAN-generated images with text prompts. In addition, Kim et al. [26] and Avrahami et al. [40] propose ways to utilize the diffusion model guided by CLIP models for global text-to-image synthesis, as well as local image editing. Text-driven generative models work well for objective things with clear “word” prompts, such as ‘apple’, ‘banana’, and ‘orange’, whereas it is barely possible to find accurate prompts for abstract image styles and image visual qualities. Moreover, prompt engineering for abstract things requires domain experts to annotate each image, and the more challenging is the similar prompts may get different CLIP score [28] (see Fig. 1). This limits the user to custom domain generation and manipulation for the abstract semantic concepts with the pre-trained CLIP.
To overcome the problems above, fine-tuning the CLIP model with considerable data for domain adaptation is a recent emerging research [33], but it is labor-intensive and time-consuming. Concurrently, prompt learning with frozen CLIP model is the alternation to extract accurate low-level image representations [28], such as luminance, exposure, and contrast [27]. We also use prompt learning to train a CLIP-Classifier, distinguishing between in-air natural and underwater degraded images. Then, the gradients of this CLIP-Classifier are used to guide the sampling process of the diffusion model.
Figure 2: The preparation for the pre-trained model. (a) Randomly select template A from the template pool (underwater domain). Then, the Color Transfer module, guided by template A, degrades a in-air natural image from INaturalist [34] into underwater domain, constructing paired datasets for training image-to-image diffusion model. (b) The image-to-image diffusion model SR3 [25] is trained to learn the prior knowledge, the mapping from the real underwater degradation domain and the real in-air natural domain, and to generate the corresponding enhancement results under the condition of the input synthetic underwater images produced by Color Transfer.
III Proposed Method
III-A Preliminary of Diffusion Models
Diffusion probabilistic models are a type of latent variable models that define a diffusion Markov chain of diffusion process to slowly inject random noise to the data, and a corresponding reverse process to slowly remove the noise [22]. The diffusion process is a forward process that gradually injects noise into the data sample ${x}_{0}$ over $T$ time steps, diffusing the data samples through Gaussian transition:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I}),$ (1)
where $\left\{\beta_{t}\right\}_{t=1}^{T}$ is an increasing variance schedule with values of $(0,1)$. The intermediate variable ${x}_{t}$ can be expressed as:
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\epsilon\sim\mathcal{N}(0,\mathbf{I}),$ (2)
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}={\textstyle\prod_{i=1}^{t}\alpha_{i}}$. The final diffuse distribution sample ${x}_{T}$ is prior random Gaussian noise.
To recover the data from the random noise ${x}_{T}$ mentioned above, the reverse process involves iterative denoising [22], which is parametrized by another Gaussian transition:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{\theta}(x_{t},t)\mathbf{I}),$ (3)
where $\mu_{\theta}(x_{t},t)$ is a function of the mean, and $\sigma_{\theta}(x_{t},t)$ is a fuction of the variance. Some works [22, 45] set the variance of the two Gaussians of the forward and reverse processes to match exactly, and the diffusion model just need to optimize $\mu_{\theta}(x_{t},t)$. Then, when $x_{t}$ is available, $\mu_{\theta}(x_{t},t)$ can be reparameterized as:
$\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)),$ (4)
where $\epsilon_{\theta}(x_{t},t)$ is a neural network resembles a function approximator trained to predict $\epsilon$ from $x_{t}$, and and the optimizing objective of Evidence Lower Bound (ELBO) is converted from $\mu_{\theta}(x_{t},t)$ to $\epsilon_{\theta}(x_{t},t)$, as follows:
$\mathcal{L}_{1}=\min_{\theta}\mathbb{E}_{x_{0}\sim q(x_{0}),\epsilon\sim\mathcal{N}(0,\mathbf{I}),t}\left\|\epsilon-\epsilon_{\theta}(x_{t},t)\right\|^{2}.$ (5)
For another understanding for the diffusion model, Song et al. [31, 46] unifies the diffusion model, a discrete multi-step denoising process, into a special form of a continuous stochastic differential equation (SDE), and the reverse process to simulate the score of each marginal distribution $\bigtriangledown\log p(x_{t})$ from $T$ to 00. Recall that $\bigtriangledown$ is shorthand for $\bigtriangledown_{x_{t}}$ in the interest of brevity. Meanwhile, the score matching of the score-based generative model is equivalent to ELBO optimization, and the score $\bigtriangledown\log p(x_{t})$ can be represented as the linear mapping of $\epsilon_{\theta}(x_{t},t)$ with the following relationship:
$\bigtriangledown\log p(x_{t})=-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t).$ (6)
For conditional generation models, Dhariwal et al. [24] uses gradients from a classifier $p(y|x_{t})$ to guide the reverse process, and by Bayes rule, the score function of this is formed as
$\bigtriangledown\log p(x_{t}|y)$ $=\bigtriangledown\log(\frac{p(x_{t})p(y|x_{t})}{p(y)})$ (7)
=$\bigtriangledown$= ▽ $\log p(x_{t})+\bigtriangledown\log p(y|x_{t})-\bigtriangledown\log p(y)$
=$\bigtriangledown$= ▽ $\log p(x_{t})+\bigtriangledown\log p(y|x_{t}).$
Notably, the gradient of $\bigtriangledown\log p(y)$ with respect to $x_{t}$ is zero, $\bigtriangledown\log p(x_{t})$ is defined as unconditional score, and $\bigtriangledown\log p(y|x_{t})$ is defined as adversarial gradient [45]. Then, $\epsilon_{\theta}(x_{t},t)$ is redefined as $\epsilon_{\theta}(x_{t},y,t)$ which corresponds to the score of the joint distribution [24]:
$\epsilon_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},t)-\sqrt{1-\bar{\alpha}_{t}}\bigtriangledown\log p(y|x_{t}).$ (8)
Simultaneously, the optimizing objective is also rewritten as:
$\mathcal{L}_{2}=\min_{\theta}\mathbb{E}_{x_{0}\sim q(x_{0}),\epsilon\sim\mathcal{N}(0,\mathbf{I}),y,t}\left\|\epsilon-\epsilon_{\theta}(x_{t},y,t)\right\|^{2}.$ (9)
After training, the sampling procedure for regular DDPM [22] or DDIM [23] can be used. It is found that Eq. (7) provides a theoretical basis for multi-guidance when conditions are independent of each other.
III-B Pipeline of Underwater Enhancement
III-B1 Multi-Classifier for Diffusion Models
Our method leverages a diffusion model for underwater image enhancement, and we use the image-to-image model SR3 [25] as the base diffusion framework. To condition the SR3 on the input $y$, the input $y$ is concatenated with $x_{t}$ along the channel dimension (see in Fig. 2 (b)), and experiments have shown that simple concatenation yields similar generation quality and makes the model more robust [25].
For conditional diffusion models, the most popular form of guidance is known as Classifier Guidance [32, 31], which is derived from score-based generation models. Our goal is to train a neural network to predict $\bigtriangledown\log p(x_{t}|y)$ the score of the conditional model, at arbitrary transition steps. Recall the score function from Equation (7):
$\bigtriangledown\log p(x_{t}|y)=\bigtriangledown\log p(x_{t})+\bigtriangledown\log p(y|x_{t}),$
where $y$ is an underwater image in this work, and of course, it can be other conditions like prompts, Canny edges, etc [21]. Following Song et al. [31] and Luo et al. [45], the Classifier Guidance can be interpreted as learning an unconditional score function combined with the adversarial gradient of a classifier $p(y|x_{t})$, and building on this observation, we will later expand the single classifier into multi-classifiers. In the following, we divide our approach into two stages: the stage-one is based on the single classifier to learn prior knowledge of the synthetic paired dataset, and the stage-two is based on the multi-classifier to fine-tune the model with additional dataset.
Stage-one: Single Classifier Guidance. Given the paired dataset synthesized by color transfer, denoted $\mathcal{D}=\left\{(x_{i},y_{i})\right\}_{i=1}^{N}$, and we are interesting in training the conditional diffusion model to learn the prior knowledge mapping from the real underwater degradation domain to the real in-air natural domain with the learning objective $\mathcal{L}_{2}$ (refer to Eq. (9)), over $x$ is an in-air natural image, where the condition $y$ is a synthesize degraded image (see in Fig. 2 (b)). Utilizing this prior knowledge can counteract the adverse impact of unreal synthetic. However, this model will be guided by the latter condition without real-underwater reference, and the fine-tuning, to directly continue training with the additional paired data, may destroy the prior knowledge.
To avoid this issue, thanks to the Classifier Guidance strategy, we first train a classifier with prior knowledge, distinguishing between in-air natural domain images and underwater domain images. Then, we exploit this classifier and the additional paired datasets to jointly fine-tune the diffusion model. As shown in Fig. 3, the enhancement results are guided by the combination of the reference domain and the in-air natural domain.
Stage-two: Multiple Classifiers Guidance. Using Bayes rule, the score function of the multi-condition model under Classifier Guidance can be derived as:
$\bigtriangledown\log p(x_{t}|y_{1},y_{2})$ $=\bigtriangledown\log(\frac{p(x_{t})p(y_{1},y_{2}|x_{t})}{p(y_{1},y_{2})})$ (10)
=$\bigtriangledown$= ▽ $\log p(x_{t})+\bigtriangledown\log p(y_{1},y_{2}|x_{t}).$
We assume that the conditions are independent of each other, and further derive the following formula:
$\bigtriangledown\log p(y_{1},y_{2}|x_{t})=\bigtriangledown\log p(y_{1}|x_{t})+\bigtriangledown\log p(y_{2}|x_{t}).$ (11)
Then, substituting this into Equation (10), we have:
$\bigtriangledown\log p(x_{t}|y_{1},y_{2})=$ $\bigtriangledown\log p(x_{t})+\bigtriangledown\log p(y_{1}|x_{t})$ (12)
$+\bigtriangledown\log p(y_{2}|x_{t}),$
where $p(y_{1}|x_{t})$ is the same as $p(y|x_{t})$ in Eq. (7), and $p(y_{2}|x_{t})$ is the trained classifier. By simply recursion, it is possible to achieve multi-conditions guidance by simple adding conditional independent classifiers.
To introduce fine-grained control to either encourage or discourage the model to move in the direction close to the in-air natural domain, we use a hyperparameter term $\lambda\in[0,1]$ to scale the adversarial gradient of the classifiers as follows:
$\bigtriangledown\log p(x_{t}|y_{1},y_{2})=$ $\bigtriangledown\log p(x_{t})+\lambda\bigtriangledown\log p(y_{1}|x_{t})$ (13)
$+(1-\lambda)\bigtriangledown\log p(y_{2}|x_{t}),$
when $\lambda<0.5$, the conditional diffusion model prioritizes the conditioning information $y_{2}$ and moves more towards the in-air natural domain. The value of $\lambda$ is empirically set to 0.4 in our experiments. Meanwhile, according to Eq. (6) and Eq. (9), we rewrite $\epsilon_{\theta}(x_{t},t)$ as $\epsilon_{\theta}(x_{t},y_{1},y_{2},t)$, combining the adversarial gradient of the two classifiers:
$\epsilon_{\theta}(x_{t},y_{1},y_{2},t)=$ $\epsilon_{\theta}(x_{t},t)-\lambda\sqrt{1-\bar{\alpha}_{t}}\bigtriangledown\log p(y_{1}|x_{t})$ (14)
$-(1-\lambda)\sqrt{1-\bar{\alpha}_{t}}\bigtriangledown\log p(y_{2}|x_{t}),$
and the overall learning objective $\mathcal{L}_{3}$ of the multi-classifiers diffusion model is depicted in Equation (15):
$\mathcal{L}_{3}=\min_{\theta}\mathbb{E}_{x_{0}\sim q(x_{0}),\epsilon\sim\mathcal{N}(0,\mathbf{I}),y_{1},y_{2},t}\left\|\epsilon-\epsilon_{\theta}(x_{t},y_{1},y_{2},t)\right\|^{2},$ (15)
where this learning objective $\mathcal{L}_{3}$ is directly used in fine-tuning the pre-trained diffusion model with additional paired datasets.
Figure 3: New fine-tuning strategy. From state $x_{T}$, according to Eq. 9, we first adopt a single classifier guidance strategy, using condition $y_{1}$—the input source image—to guide the reverse diffusion process until state $x_{t}$. Then, we switch to the multi-classifier guidance strategy according to Eq. 15. With multi-condition guidance, the intermediate results from $x_{t}$ to $x_{0}$ are constrained by both the natural and reference domains, mitigating the damage of fine-tuning to the prior knowledge of the pre-trained model.
III-B2 Color Transfer for Degraded Image Synthesis
The color transfer can realize the translation between image domains, by transfering one image’s color characteristics to another [29]. Naturally, we can use this method to translate the image from the in-air natural domain into the underwater scene domain.
Recall that the color transfer align the color appearance of a source image $I_{s}$ with a target image $I_{t}$ (all images have been converted from RGB to CIELAB color space ($L^{\ast}a^{\ast}b^{\ast}$)):
$l_{s}^{\prime}$ $=\frac{\sigma_{t}^{l}}{\sigma_{s}^{l}}(l_{s}-\left\langle I_{s}^{l}\right\rangle)+\left\langle I_{t}^{l}\right\rangle,$ (16)
$a_{s}^{\prime}$ $=\frac{\sigma_{t}^{a}}{\sigma_{s}^{a}}(a_{s}-\left\langle I_{s}^{a}\right\rangle)+\left\langle I_{t}^{a}\right\rangle,$
$b_{s}^{\prime}$ $=\frac{\sigma_{t}^{b}}{\sigma_{s}^{b}}(b_{s}-\left\langle I_{s}^{b}\right\rangle)+\left\langle I_{t}^{b}\right\rangle,$
[…]
In our case, for each in-air natural image $I_{s}$, we randomly select one image as Template A from Template Pool (the underwater scene domain) as the target image $I_{t}$, and then use $I_{t}$ to guide $I_{s}$ to transfer to the underwater scene domain via Equation (16). The detail process is depicted in Fig. 2 (a). Finally, we convert the degraded synthetic image back to RGB color space.
Figure 4: Illustration of the prompt learning for CLIP-Classifier. (A) Prompt Initialization. Given two text prompts describing the in-air natural image $I_{n}$ and underwater image $I_{u}$. We encode each text and get the initial in-air natural image prompt $T_{n}\in\mathbb{R}^{N\times 512}$ and the initial underwater image prompt $T_{u}\in\mathbb{R}^{N\times 512}$. (B) Prompt Training. We use the cross-entropy loss to constrain the learnable prompts, aligning learnable prompts and images in the CLIP latent space, by maximizing the cosine similarity for matched pairs $\left\{T_{n},I_{n}\right\}$ and $\left\{T_{u},I_{u}\right\}$. The base model of CLIP is frozen throughout. The learned prompt for CLIP-Classifier will be used in Fig. 3
III-B3 Prompt Learning for CLIP-Classifier
In Section. III-B1, what we want is a classifier to control the diffusion model generation process, making the result move toward the in-air natural domain. Thanks to the CLIP’s near-perfect text-driven property [43], we design a CLIP-Classifier to implement this vision, and all we need to do is to find prompts that can distinguish between in-air natural images and underwater images. However, prompts for the quality and abstract perception of natural and underwater images often require a lot of manual prompt tuning, and the similar prompts easily lead to huge differences in CLIP score. To overcome this issue, CoOp [28] models a prompt’s context words with learnable vectors to broader CLIP-like vision-language models to downstream applications. For our task, we only need two prompts to characterize in-air natural and underwater images. Therefore, similar to CLIP-LIT [27], we utilize the prompt learning to model a prompt’s the longest text with learnable tensors, and maximize the ability to characterize images. Note that the base model of CLIP is frozen in the entire process. The process of prompt learning involves two key components: prompt initialization and prompt training, as detailed below.
Prompt Initialization. Randomly given two texts describing in-air natural and underwater images respectively, CLIP converts each text into the multi-model embedding space, obtaining the in-air natural image prompt $T_{n}\in\mathbb{R}^{N\times 512}$, and the underwater image prompt $T_{u}\in\mathbb{R}^{N\times 512}$ (refer to Fig. 4 (a)). $N$ represents the length of prompt tokens, with an upper limit of 77. CoOp [28] has experimentally shown that more prompt tokens lead to better performance. Thus, we model $T_{n}$ and $T_{u}$ as learnable tensors with $N=77$, maximizing the length of prompt tokens.
Prompt Training. Given an in-air natural image $I_{n}$ and an underwater image $I_{u}$, the image encoder $\Phi(\cdot)$ of the frozen CLIP takes them as input and extracts image features. On the other hand, by forwarding the prompt $T_{n}$ and the underwater prompt $T_{u}$ to the text encoder $\Psi(\cdot)$, we can obtain the text features. Given a batch of image-text pairs, the prediction probability is calculated as follows:
$P(T_{i}\mid I)=\frac{e^{cos(\Psi(T_{i}),\Phi(I))}}{\sum_{i\in\left\{n,u\right\}}e^{cos(\Psi(T_{i}),\Phi(I))}},$ (17)
where $I$ is an image, $I\in\left\{I_{n},I_{u}\right\}$ and $cos(\cdot,\cdot)$ denotes cosine similarity. Specially, the $T_{n}/T_{u}$ is shared among all in-air natural/underwater images. Similar to CLIP [43], the prompt learning is trained to align learnable prompts and images in embedding spaces, by maximizing the cosine similarity for matched pairs $\left\{T_{n},I_{n}\right\}$ and $\left\{T_{u},I_{u}\right\}$, while minimizing the cosine similarity for incorrect pairs. The cross-entropy loss can be used as a learning objective, and its gradients can be propagated all the way back to update the prompt tensors (see in Fig. 4 (b)). Since there are only two prompts in our prompt engineering, we can simplify the process by directly using the binary cross-entropy to classify in-air natural and underwater images to learn the prompt tensors:
$\mathcal{L}_{p}=-(q\ast logP(T_{n}\mid I)+(1-q)\ast log(1-P(T_{n}\mid I))),$ (18)
Where $q$ is the label of one-hot, which is $1$ for in-air natural images and 00 for underwater images.
CLIP-Classifier. After training, we can get the well-performing prompts and then construct the CLIP-classifier by measuring the similarity between the generation image and the prompts in the CLIP space:
$\mathcal{L}_{clip}=\frac{e^{cos(\Psi(T_{u}),\Phi(I_{g}))}}{\sum_{i\in\left\{n,u\right\}}e^{cos(\Psi(T_{i}),\Phi(I_{g}))}},$ (19)
Where $I_{g}$ is the image generated by the diffusion model.
Substituting CLIP-Classifier $\mathcal{L}_{clip}$ into Equation (14), the new $\epsilon_{\theta}(x_{t},y_{1},y_{2},t)$ is:
$\epsilon_{\theta}(x_{t},y_{1},y_{2},t)$ $=\epsilon_{\theta}(x_{t},t)-\lambda\sqrt{1-\bar{\alpha}_{t}}\bigtriangledown\log p(y_{1}|x_{t})$ (20)
$-(1-\lambda)\sqrt{1-\bar{\alpha}_{t}}\bigtriangledown\log\mathcal{L}_{clip}(y_{2},x_{t}),$
Where due to the properties of the cosine function, $y_{2}=T_{u}$. Next, we explore the details of the CLIP-Classifier $\mathcal{L}_{clip}$ guides generation process.
Figure 5: Qualitative analysis of the effectiveness of the CLIP-Classifier. (a) CLIP score of the intermediate variable $x_{t}$ of the image versus time $t$. The time step is set to 2000. The CLIP score curves of the raw and reference images are plotted separately. (b) We calculate the difference in CLIP score between the raw and reference images, then plot the difference curve and label several key time points on the curve.
III-C Fine-Tuning Strategy
The forward process of the diffusion model is actually a process in which image semantic information is continuously obscured by noise, always starting from high-frequency information to low-frequency information. In contrast, the reverse generation process smoothly converts an initial Gaussian noise to a realistic image sample through iterative denoising, working from low-frequency information to high-frequency information. For special image-to-image tasks that require consistent content, such as super-resolution [25], image restoration [47], repaint [48], etc., the input image and the corresponding image share low-frequency content in each transition from $x_{t}$ to $x_{t-1}$ during the generation process [49]. Therefore, image synthesis and editing on diffusion model can start from the high-frequency region. For instance, SDEdit [50] is to “hijack” the generation process, adding a suitable amount of noise to smooth out high-frequency semantic details, while still preserving the overall structure of the input image, and then initialize the SDE with this noisy input, obtaining a denoised result that is faithful to the user guidance input.
Obviously, in our case, the underwater degradation images and the reference images share low-frequency content, so we can introduce new conditions at intermediate moments $x_{t}$ to guide the diffusion process (see in Fig. 3). In addition, given a batch of data pairs, we first use Eq. (2) to produce a sequence of intermediate variables $(x_{0},x_{1},\dots,x_{T})$ for each image, by perturbing images with Gaussian noise at time $t\in(0,1)$, and then employ CLIP-Classifier to calculate the classification probability (CLIP score), as illustrated in Fig. 5 (a). Notably, the randomness of the perturbation noise directly causes the CLIP score curves to vibrate violently in small intervals, but the overall trend of the curves is still pronounced. From Fig. 5 (b), we observe that the CLIP-Classifier $\mathcal{L}_{clip}$ mainly works in the discontinuous high-frequency region where $t\in(0,0.10)\cup(0.25,0.50)$, and other range can not be distinguished by the CLIP-Classifier. Combined with the diffusion model generative prior, sharing the low-frequency content, we only need to guide the pre-trained diffusion model in the high-frequency region during the fine-tuning process. The new strategy improves the fine-tuning speed compared to the traditional strategy, and can even speed up to 10 times, as will be discussed in detail in Section IV-F.
IV Experiments
IV-A Benchmarks and Evaluation Metrics
Prior Knowledge. It is well known that real reference images are difficult to obtain for underwater image enhancement tasks. We adopt the synthetic dataset strategy, using color transfer to degrade the in-air natural images of INaturalist 2021 (mini) [34] to the underwater domain to yield paired datasets (see in Fig. 2 (a)). INaturalist 2021 (mini) has 10,000 categories with 50 images in each category, totalling 500k images. The image-to-image diffusion model is trained from scratch on this dataset to learn prior knowledge from the real underwater degradation domain to the in-air natural domain.
Dataset. The direct finetuning or continued training of a large pretrained image-to-image diffusion model with limited data may cause overfitting and catastrophic forgetting [21]. The SUIM-E [11] has only 1635 images, while UIEB [17] has only 890 images, which is significantly smaller than the INaturalist dataset used to pre-train the diffusion model. Therefore, we utilize the pre-divided SUIM-E-Train set, which consists of 1525 paired underwater images, and 800 paired images randomly selected from UIEB as the training set, for fine-tuning our model. The test set from SUIM-E and the remaining 90 images from UIEB are combined to form a test set of 200 images, called T200.
We also conduce tests on the Challenging set of UIEB, for which contains 60 tough cases for enhancement without references, called C60. In addition, we further verify the generalisation and robustness of our model on the following commonly used underwater image datasets, which do not have reference images. SQUID [51] contains 57 images taken in different locations, enabling a rigorous evaluation of restoration algorithms on in-air natural images. Color-Checker7 [52] covers 7 underwater images taken with the camera from different angles in a shallow swimming pool, which is also used to evaluate the accuracy of the proposed CLIP-UIE for color correction.
Evaluation Metrics. Since not all datasets have reference images, we divide five commonly used evaluation metrics into two groups for comprehensive quantitative evaluation. For full-reference evaluation metrics, we use two metrics: Peak Signal to Noise Ratio (PSNR) which measures the ratio of the peak signal to the average energy, and Structural Similarity (SSIM[53] which compares the target and ground-truth images in terms of brightness, contrast, and structure. For non-reference evaluation metrics, we utilize three metrics: Underwater Image Quality Measure (UIQM[54] which considers the underwater contrast, colorfulness, and sharpness on HSV space, Underwater Color Image Quality Evaluation (UCIQE[55] which is based on chroma, contrast, and saturation to quantify the nonuniform color cast, blurring, and low-contrast for the degraded underwater images, and Cumulative Probability of Blur Detection (CPBD[56] which utilizes a probabilistic model to estimate the probability of detecting blur at each edge in the image.
TABLE I: Quantitative comparison on T200, C60, SQUID and Color-Check7 datasets. Two reference-based metrics (i.e., PSNR and SSIM) and three Non-reference quality metrics (i.e., UIQM, UCIQE and CPBD) are adopted. Double horizontal lines separate the traditional and deep learning-based approaches. The best scores and second-best scores are marked in Red and Blue, respectively. (Best viewed in color)
Datasets T200 C60 SQUID Color-Checker7
[…]
UDCP 11.803 0.548 1.268 0.598 0.636 0.952 0.552 0.551 0.680 0.575 0.627 1.642 0.641 0.589
ULAP 16.570 0.768 0.955 0.614 0.647 0.690 0.579 0.560 0.466 0.511 0.659 0.752 0.628 0.557
Ucolor 21.907 0.888 0.586 0.587 0.606 0.505 0.553 0.512 0.182 0.522 0.637 0.693 0.586 0.580
TCTL-Net 22.403 0.897 0.796 0.608 0.617 0.576 0.587 0.510 0.133 0.547 0.656 1.092 0.613 0.583
UIEC^2-Net 23.347 0.860 0.822 0.610 0.631 0.661 0.583 0.549 0.323 0.571 0.641 0.893 0.619 0.617
UDAfomer 25.350 0.921 0.759 0.596 0.606 0.543 0.561 0.502 0.187 0.545 0.607 0.949 0.609 0.570
DM_underwater 25.569 0.931 0.797 0.609 0.576 0.647 0.579 0.492 0.242 0.560 0.612 0.874 0.614 0.576
CLIP-UIE 25.412 0.936 0.981 0.619 0.624 0.754 0.588 0.497 0.424 0.575 0.629 1.257 0.645 0.544
Figure 6: Visual comparisons on underwater images from T200 dataset. From left to right are raw underwater images and the results of UDCP [8], ULAP [10], Ucolor [1], TCTL-Net [30], UIEC^2-Net [12], UDAformer [15], DM_underwater [41], the proposed CLIP-UIE, and reference images are presented, respectively.
Figure 7: Visual comparisons on challenging underwater images from C60 dataset. From left to right are raw underwater images and the results of UDCP [8], ULAP [10], Ucolor [1], TCTL-Net [30], UIEC^2-Net [12], UDAformer [15], DM_underwater [41] and the proposed CLIP-UIE are presented, respectively.
Figure 8: Visual comparisons on challenging underwater images from SQUID dataset. From left to right are raw underwater images and the results of UDCP [8], ULAP [10], Ucolor [1], TCTL-Net [30], UIEC^2-Net [12], UDAformer [15], DM_underwater [41] and the proposed CLIP-UIE are presented, respectively.
Figure 9: Visual comparisons on underwater images from Color-Checker7 dataset. From left to right are raw underwater images and the results of UDCP [8], ULAP [10], Ucolor [1], TCTL-Net [30], UIEC^2-Net [12], UDAformer [15], DM_underwater [41] and the proposed CLIP-UIE are presented, respectively. The standard color card is appended to the upper left corner of the first column.
IV-B Implementation Details
We implement our method with the PyTorch platform. For training, we initialize all the weights of each layer with Xavier initialization [57] and employ the Adam optimizer with a learning rate of 3e-6 with the scheduler of linear decay to zero. Moreover, batch size is set to 8, and the time step of the diffusion model is set to 2000 (namely T=2000). To ensure a fair comparison, images for all compared methods were resized to 256×256 during both training and testing. Additionally, our data augmentation techniques include random rotation and random horizontal flips. The experiments are conducted on a consistent hardware setup comprising a single NVIDIA RTX 3090 GPU, a 2.3GHz Intel Xeon processor, 128GB RAM, and the Ubuntu 18.04 operating system platform. The pre-trained model we used has been trained from scratch for 300 hours on INaturalist 2021 (mini).
IV-C Compared Methods
We compare our method CLIP-UIE with seven state-of-the-art UIE approaches, including two typical traditional methods (UDCP [8] and ULAP [10]), and five deep learning-based methods (Ucolor [1], TCTL-Net [30], UIEC^2-Net [12], UDAformer [15], DM_underwater [41]). The aforementioned two UIE methods are the most representative physical model-free methods that have been widely adopted in recent years. The deep learning-based approaches compared here have achieved state-of-the-art (SOTA) performance over the past years. Among them, the Ucolor [1] is based on medium transmission-guided multi-color space embedding, the TCTL-Net [30] is based on template-free color transfer, UIEC^2-Net [12] is based on RGB and HSV Color Space, UDAformer [15] employs dual attention transformer, and DM_underwater [41] is the first UIE method based on diffusion model. To ensure a fair and rigorous comparison, we utilized the same training set from the SUIM-E and UIEB dataset to fine-tune or continued train all the deep learning-based methods. Since the Ucolor [1] training phase requires special image sizes, we only use 256x256 image sizes for the test phase. Other comparison methods use 256x256 image size in both training and testing phases. In addition, all the comparison methods utilize the source codes released by the respective authors to produce their results and strictly adhere to the identical experimental settings across all evaluations.
TABLE II: The color dissimilarity comparisons of different methods on color-check7 in terms of the CIEDE2000. The smaller the value, the better the performance. The Top Three Scores are in Red, Blue, Green. (Best Viewed in Color)
Method D10 Z33 T6000 T8000 TS1 W60 W80
Input 12.91 16.65 14.99 19.30 16.15 11.97 14.12
UDCP 25.76 24.54 21.79 32.99 29.04 23.57 24.54
ULAP 26.46 17.79 18.25 25.01 19.26 24.75 18.36
Ucolor 11.66 15.11 17.24 18.59 15.93 10.15 13.58
TCTL-Net 12.22 14.96 13.51 20.33 16.34 10.70 16.64
UIEC^2-Net 17.81 21.86 17.55 19.61 20.82 16.11 11.59
UDAformer 11.99 16.15 13.98 21.62 14.11 12.09 11.10
DM_underwater 13.19 13.97 16.78 19.68 17.89 13.64 12.09
CLIP-UIE 10.18 14.43 12.10 20.36 15.52 11.56 12.39
IV-D Comparison on Visual Quality of Enhancement
The quantitative comparisons on the T200, C60, SQUID and Color-Check7 datasets are reported in Table I. We mainly use PSNR and SSIM as our quantitative indices for full-reference T200 datasets, and UIQM, UCIQE and CPBD for non-reference C60, SQUID and Color-Checker7 datasets. The results in Tab. I show that our CLIP-UIE outperforms the other compared methods, and achieve state-of-the-art performance in terms of image quality evaluation metrics on UIE task, particularly in UIQM and UCIQE metrics, with an improvement of 0.1 and 0.01, respectively. This verifies that our method can produce more natural enhanced images, which better corrects degraded colors and maps from underwater degraded domains to in-air natural domains. To better validate the superiority of our methods, Figure 6 shows the visual results of different comparison methods on T200 dataset. The traditional methods like UDCP and ULAP, usually cause color distortions and damage image detail (see the examples in Fig. 6 (b, c)), albeit with high non-reference metrics, which also occurs in other datasets. The deep learning-based methods can produce pleasing visual enhancement due to the direct use of reference images as constraints during the training phrase. However, this limits the enhanced image to a human-selected reference domain, and some methods have very similar enhancement results and do not outperform the reference (e.g. (d - h) in the second and fourth rows of Fig. 6). On the contrary, our CLIP-UIE introduces the prior knowledge of in-air natural domain through CLIP-Classifier, which can produce more natural images, breaking through the limitation imposed by reference domain to a certain extent, as shown in Fig. 6 (i).
To further verify the effectiveness and robustness of our method, we conducted experiments on C60, SQUID and Color-Checker7 datasets. (1) The visual comparisons with other methods on the C60 dataset are shown in the Fig. 7. The underwater images suffer from lower contrast and more severe color casts (see the examples in the first row of Fig. 7), which raise great challenges for the enhancement methods. By observation, the deep learning-based methods show better robustness and generalisation than the traditional methods. As a comparison, our method has better performance in color correction and detail restoration. Specially, our method improves 11% in UIQM metrics. (2) On the SQUID dataset, our CLIP-UIE outperforms all other deep learning-based methods and achieves optimal scores in terms of underwater image quality non-reference evaluation metrics including UIQM and UCIQE, with 0.424 and 0.575, respectively. (see in Table I). As shown in Figure 8, our method performs more natural visual enhancement for underwater images taken at different depths, which verifies the robustness of the proposed CLIP-UIE. This phenomenon can be attributed to the conditional generation is jointly controlled by the source image $y_{1}$ and CLIP-Classifier $y_{2}$ (the process is shown in Fig. 3), thereby the multi-conditions guidance enhances generation quality. (3) We further conduce comparative tests on the Color-Checker7 dataset to evaluate the generalization and robustness of our model. Figure 9 visually illustrates the difference between output images of CLIP-UIE and other methods. In the first row of Fig. 9, the difference lies in the correction of skin color, where the image generated by CLIP-UIE is closer to the real colors. In the second row of Fig. 9, the ULAP [10] produce the enhanced image with good background lighting, but there is still a color deviation due to the introduction of excessive red components. Comparing to deep learning-based methods, the results produced by our method are brighter and more visually appealing.
TABLE III: Quantitative results of ablation study on T200 datasets. The best scores are in Red. (Best viewed in color)
[…]
-w/o-CLIP 26.174 0.943 0.783 0.585 0.607
CLIP-UIE 25.412 0.936 0.981 0.619 0.624
IV-E Comparison on Color Correction Performance
In order to further illustrate the performance of the proposed CLIP-UIE for color restoration of underwater images, we conduced detailed color dissimilarity comparisons on the Color-Checker7 [52] dataset, which is photographed in a swimming pool with divers holding standard color cards. Following Ancuti et al. [10], we use CIEDE2000 [52] to measure the relative differences between the corresponding color patches of the Macbeth color checker [58] and the enhancement results of these comparison state-of-the-art methods, making it easy to understand that a standard color card is appended to the upper left corner of the first column of Fig. 9. The results, as presented in Tab. II, indicates that improve the performance of UIE methods in recovering ground-truth colors, but not very noticeably, and the enhancement result of the image T8000 is even worse than the input. We believe that the image encoder $\Phi(\cdot)$ compresses images into high-dimensional features, and CLIP-Classifier pays more attention to global features during generation process. As shown in the first row of Fig. 9, the result of our method presents better brightness and visual perception, but does not achieve a significant improvement in color correction.
Figure 10: The distribution of similarity score (CLIP score) between the learned in-air natural image prompt and images across the T200 dataset. The point-box plot and kernel density estimation curve of the reference images and the enhancements with the CLIP-UIE-w/o-CLIP and complete CLIP-UIE are presented. The closer the point-box plot is to the left, the closer it is to the in-air natural domain.
Figure 11: Ablation study of the effectiveness of CLIP-Classifier. From left to right are raw underwater images, reference images, the enhancements with the CLIP-UIE-w/o-CLIP and complete CLIP-UIE are presented, respectively. With CLIP-Classifier, CLIP-UIE breaks through the limitations of the reference domain and achieves more natural enhancement results.
IV-F Ablation Study
Effectiveness of CLIP-Classifier. To Investigate the impact of the CLIP-Classifier on our method, we fine-tune our conditional diffusion model on the same training data without CLIP-Classifier (w/o-CLIP). The scores of the main evaluation metrics on T200 datasets are given in Tab. III. The results indicate that CLIP-Classifier guides the diffusion model toward the in-air natural domain, leading to a decrease in the full-reference metrics and an increase in the non-reference metrics. Besides, we use CLIP-Classifier to evaluate the similarity (CLIP score) between the in-air natural domain and enhanced results, and the corresponding point-box plot and kernel density estimation (KDE) curve are shown in Fig. 10. The distributions of the w/o-CLIP and the Reference are very similar (see KDE in Fig. 10), which indicates that its enhancement performance is constrained by the reference domain. On the contrary, due to the introduction of the prior knowledge of in-air natural domain through CLIP-Classifier, our CLIP-UIE breaks through this limitation imposed by reference domain to a certain extent, keeping the distribution of the CLIP-UIE far away from the reference domain and close to the in-air natural domain and reducing subjective preferences.
Visual comparison illustrated in Figure 11 can also significantly support the above points. Compared with its ablated versions, the proposed CLIP-UIE achieves more natural enhancement results, with the guidance of CLIP-Classifier that inherits prior knowledge of the in-air natural domain. And without real-world underwater reference images, the combination of the reference and in-air natural domains drives the enhancement results closer to real-world natural images (see Fig. 3 and Fig. 11 (c, d)).
Figure 12: The visual comparisons of the enhancements with different ablated models and the full CLIP-UIE model on the T200 and SQUID dataset. From the left to right, (a) the raw images, the enhancements with ablated models (b) the model $t_{m}$=0.10, (c) the model $t_{m}$=0.25, (d) the model $t_{m}$=0.50, (e) the model $t_{m}$=0.75, (f) the model $t_{m}$=1.00 (the full CLIP-UIE model) and (g) the reference images are presented, respectively. In the second line, the reference image of the SQUID dataset is not available.
TABLE IV: Efficiency analysis and comparison of new fine-tuning strategy on T200 and SQUID datasets. The best and second-best scores are in Red and blue, respectively. (Best viewed in color)
Dataset T200 SQUID
[…]
$t_{m}$=0.10 24.339 0.930 0.943 0.611 0.619 0.417 0.550 0.637
$t_{m}$=0.25 23.887 0.923 0.894 0.602 0.618 0.291 0.541 0.622
$t_{m}$=0.50 24.470 0.912 0.936 0.622 0.651 0.376 0.583 0.631
$t_{m}$=0.75 25.354 0.946 0.924 0.608 0.625 0.389 0.559 0.628
$t_{m}$=1.00 25.412 0.936 0.981 0.619 0.624 0.424 0.575 0.629
Fine-Tuning Strategy. In addition to the observation provided in Section III-C, to further validate the new fine-tuning strategy, we provide the quantitative comparison in Tab. IV. The model $t_{m}$ represents the fact that we only fine-tune the model pretrained on INaturalist [34] in the range $t\in(0,t_{m})$ (the training time step $n\in(0,2000\times t_{m})$) (refer to Fig. 3). The fine-tuning range of the model $t_{m}$=0.25 is larger than that of the model $t_{m}$=0.10, but the performance does not exceed the latter, especially the main evaluation metrics PSNR and UIQM decreases from 24.339 and 0.943 to 23.887 and 0.894 respectively on the dataset T200. The drop in UIQM on the dataset SQUID is even more dramatic, from 0.417 to 0.291. We believe this may be due to the fact that the image-to-image diffusion model is more sensitive to high-frequency information, and that the CLIP-Classifier does not steer the generation process in the high-frequency range $t\in(0.1,0.25)$, instead interfering with the condition $y_{1}$ (refer to Fig. 3). Further comparison of the models $t_{m}$=0.50, $t_{m}$=0.75 and $t_{m}$=1.00 shows that the performance of the models is stable and the evaluation metrics are without large transition, which indicates that the CLIP-Classifier cannot distinguish the image pairs perturbed by Gaussian noise in the low frequency region $t\in(0.50,1.00)$ and has less impact on the overall generative process. This experimental result aligns with the Fig. 5 (b). Figure 12 presents the visual comparison between the proposed CLIP-UIE (the model $t_{m}$=1.00) and the other ablated versions. As shown in Fig. 12 (b), for the image-to-image special task, which is low-frequency content sharing, we are able to adapt to the downstream task by fine-tuning the model only in the range of $t\in(0,0.10)$ and the enhancement results are competitive, which is 10 times faster over the traditional strategy.
V Discussion and Future works
As illustrated in Fig. 5, The score distribution of the results produced by the proposed CLIP-UIE is not clustered enough, and some enhanced images are still in the reference domain. Additionally, the score difference curve of the CLIP-Classifier is different from the normal subjective assumption—the more noise, the lower the curve—and the effective range of the classifier is discontinuous $t\in(0,0.10)\cup(0.25,0.50)$, and the image pairs perturbed by Gaussian noise cannot be distinguished when $t\in(0.10,0.25)\cup(0.50,1.00)$. This makes it difficult to precisely control the generation process of diffusion models. In future research, it would be interesting to train an effective range-stable CLIP-Classifier in order to achieve precise control of the multi-guidance diffusion model, and to make the enhancement results more natural.
As a new fine-tuning strategy, the speed is faster than the traditional fine-tuning strategy, but the overall performance is not superior to traditional methods (refer to Tab. IV). The trade-off between speed and performance is a common challenge for all deep learning-based methods. Exploring how to strike a balance speed and performance in the fine-tuning of image-to-image diffusion models is also a promising topic. Unfortunately, this fine-tuning strategy is limited to the image-to-image diffusion model, which shares low-frequency content.
VI conclusion
In this paper, we have presented a novel framework CLIP-UIE for underwater image enhancement, which combines the pre-train diffusion model and CLIP-Classifier to conduct conditional generation. In our framework, the pre-train diffusion model has learned the prior knowledge of image enhancement from the underwater degradation domain to the real in-air natural domain. The customized CLIP-Classifier inherits the prior knowledge of the in-air natural domain, which is used to counteract catastrophic forgetting and mode collapse of the diffusion model during fine-tuning, allowing the enhancement results to break through the limitations of the reference domain to a certain extent. Simultaneously, the image-to-image enhancement task requires consistent content between the input and the output images, so that the diffusion model shares low-frequency information, which means that the range of fine-tuning is concentrated in the high-frequency region. Further analysis of CLIP-Classifier shows that it also mainly acts in the high-frequency region, similar to the diffusion model. Therefore, we propose a new fine-tuning strategy that guides the pre-trained diffusion model only in the high-frequency region, e.g., the model $t_{m}$=0.10—we only fine-tune in the range $t\in(0,0.10)$—achieves competitive enhancement results, while the speed is 10 times faster than the traditional strategy. Extensive experiments show that CLIP-UIE not only generates enhanced images of superior quality beyond the reference images but also surpasses other advanced UIE methods. In the future, we aim to introduce more prior knowledge by classifiers break through the limitations of the reference domain and further enhance visual quality.